TiMoThYnICee:

Hello everyone I'm new here and I want everyone to know this answer, like I do, Which of the following equations will have a negative solution? A.m - 4 2/3 = 7 1/2 B.m - 71/2 = 4 2/3 C.7 1/2 + m = 4 2/3 D.-7 1/2 + m = 4 2/3

TiMoThYnICee:

The answer would be C) Since the answer is m = -17/6. Cant be A) cause that is m = 73/6 Cant be B) cause that is m = 241/6 Cant be D) cause that is m = 73/6

TiMoThYnICee:

that's the answer, hope this helps